OpenStudy (anonymous):

Twyla and Tony have a 30/5 balloon mortgage for $389,900 with a rate of 4.85%. How much will they pay in interest over the life of the loan?

OpenStudy (anonymous):

I just need help trying to figure out how someone would solve this...

OpenStudy (amistre64):

a 30 year mortgage that is paid on if 5 years, then the balance is due

OpenStudy (amistre64):

there is a formula i tend to use for finding stuff like this\[B_n=B_o(k)^n-P~\frac{1-(k)^n}{1-(k)}\] since its a 30 year loan, and i assume its monthly payment; 30*12 = 360, n= 360 \[0=B_o(k)^{360}-P~\frac{1-(k)^{360}}{1-(k)}\] \[B_o(k)^{360}\frac{1-(k)}{1-(k)^{360}}=P\] i use k to clean up it up; but k = 1+r/12 in this case \[k=1+\frac{.0485}{12}\approx 1.00404\] \[B_o(1.00404)^{360}\frac{1-(1.00404)}{1-(1.00404)^{360}}=P\] \[B_o(1.00404)^{360}\frac{.00404}{1-(1.00404)^{360}}=P\] and for the wolf that would look like: 389900(1.00404)^(360) (.00404)/(1-(1.00404)^(360)) http://www.wolframalpha.com/input/?i=389900%281.00404%29%5E%28360%29+%281-1.00404%29%2F%281-%281.00404%29%5E%28360%29%29 so payments of 2057 a month

OpenStudy (amistre64):

so the question is, how much has been paid in 5 years, how much of the balance is left, and the workings of that should determine the interest paid

OpenStudy (anonymous):

I know how much there was monthly, because I'm sure that has some to do with the problem, I just don't know where to go from there.

OpenStudy (amistre64):

im just doing some thinking here to see how close i come \[B_{72}=B_o(k)^{72}-2057~\frac{1-(k)^{72}}{1-(k)}\]\[~~~~~=349,731\] thats how much we would have left to pay off, thats the balloon payment how much have we paid so far? 72 * 2057 = 148,104, entotal. to find how much that was in principal, i would take the difference between the loan amount and the balance 389,900 - 349,741 ----------- 40,259 ; thats the principal paid, the rest is interest 148,104 -40,259 -------- 107,845

OpenStudy (amistre64):

if you know your payment amount, how much is it?

OpenStudy (anonymous):

The only info I have is what shows in the question, and the answer options. Would it help if I attached the answer options?